OpenStudy (anonymous):

find the value of Arc cos square root of 2 over 2

OpenStudy (dangerousjesse):

I guess you can just use the inverse cosine value \(\large (cos^{-1}(x))\) to get \(\large \frac{1}{2} cos^{-1}(\sqrt 2)\).

OpenStudy (anonymous):

\[\cos ^{-1}\left(\frac{\sqrt{2}}{2}\right)=\frac{\pi }{4}=45 {}^{\circ} \]